7t(n/2)+18(n/2)^2 I need a help to know how to solve this
Answer #1

Solution is provided below. If any doubt please comment below.

The given function is T(n) = 7T(n/2) + 18(n/2)2.

T(n) = 7T(n/2) +(18/4)n2

Apply master’s theorem

Compare this equation with T(n) = aT(n/b) + f(n)

Here a=7, b=2 and f(n) = Θ(nc)= Θ (n2)

Here c= 2 and c<log27

Therefore , T(n) = Θ (nlogba) = Θ (nlog27) = Θ (n2.8)